# Descent-Guided Policy Gradient for Scalable Cooperative Multi-Agent Learning

### v1

@@ E.1 Lemma: Guidance Consistency
$=(1-\mu\eta)\Delta_{k}+\frac{L\eta^{2}}{2}\sigma_{\text{DG}}^{2}.$
Unrolling the recurrence for $T$ iterations:
$\Delta_{T}\leq(1-\mu\eta)^{T}\Delta_{0}+\frac{L\eta}{2\mu}\sigma_{\text{DG}}^{2}.$
To achieve $\Delta_{T}\leq\epsilon$, it suffices to make each term at most $\epsilon/2$.
Steady-state term. Requiring $\frac{L\eta}{2\mu}\sigma_{\text{DG}}^{2}\leq\frac{\epsilon}{2}$ gives the constraint $\eta\leq\frac{\mu\epsilon}{L\sigma_{\text{DG}}^{2}}$. We set $\eta=\frac{\mu\epsilon}{L\sigma_{\text{DG}}^{2}}$ (the largest admissible step size, which minimizes the number of iterations needed).
Transient term. Requiring $(1-\mu\eta)^{T}\Delta_{0}\leq\frac{\epsilon}{2}$ and using $\log(1-x)\leq-x$, we need $T\geq\frac{1}{\mu\eta}\log\frac{2\Delta_{0}}{\epsilon}$. Substituting $\eta=\frac{\mu\epsilon}{L\sigma_{\text{DG}}^{2}}$:
$T\geq\frac{1}{\mu\cdot\frac{\mu\epsilon}{L\sigma_{\text{DG}}^{2}}}\log\frac{2\Delta_{0}}{\epsilon}=\frac{L\sigma_{\text{DG}}^{2}}{\mu^{2}\epsilon}\log\frac{2\Delta_{0}}{\epsilon}.$
@@
By Theorem 4.2, $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ (independent of $N$), so the sample complexity is:
$T=\mathcal{O}\left(\frac{L}{\mu^{2}\epsilon}\log\frac{1}{\epsilon}\right).$
Since neither $L$, $\mu$, nor $\sigma_{\text{DG}}^{2}$ depends on the number of agents $N$, the sample complexity $T=\mathcal{O}(1/\epsilon)$ is independent of $N$, completing the proof. ∎
@@ H.1 Physical Environment
The efficiency parameters $\eta_{\text{CPU}}$ and $\eta_{\text{Mem}}$ model performance-per-watt differences across hardware generations: Gen-1 instances (circa 2015) have $\eta_{\text{CPU}}\in[0.70,0.75]$, Gen-2 (circa 2017) have $\eta_{\text{CPU}}\in[0.84,0.98]$, and Gen-3 (circa 2020+) have $\eta_{\text{CPU}}\in[1.06,1.08]$. Job completion time scales as $\text{duration}/\eta_{\text{CPU}}$, making high-efficiency servers more productive. Server types are sampled according to the specified weights and randomly shuffled across indices at scenario generation to prevent agents from exploiting positional patterns.
@@ H.3 Reward Function
where $Q_{\text{global}}(t)$ is the global buffer size, $Q_{k}(t)$ is the local queue at server $k$, $L_{k}(t)$ is the workload (CPU + memory) at server $k$, $\eta_{k}$ is the server’s efficiency, and $C_{\text{total}}$ is the total cluster capacity. We set $w_{1}=1$ and $w_{2}=20$ so that both penalties contribute comparably to the reward: the queue penalty typically ranges in $[0,20]$, whereas the normalized energy ratio remains below $1$. Load-balancing (utilization variance) is handled entirely by the DG-PG guidance signal and is therefore excluded from the reward.
@@ I.2 Linear Models (Controlled Comparison)
  • Large mini-batch (10,000) and high rollout count (24): With smaller batches (e.g., 256–1,024), MAPPO and IPPO exhibit wildly oscillating rewards that never converge. The high gradient variance of multi-agent policy gradients in this environment—heterogeneous servers, bimodal workloads, and non-stationary arrivals—requires large batches to average over noise.
@@ I.3 Neural Network Models (DG-PG)
  • Guidance weight $\alpha$: dynamic three-phase schedule (Section 6.2): $\alpha=0.9$ for the first 10% of training, linear decay $0.9\to 0.2$ from 10% to 50%, and $\alpha=0.2$ thereafter
@@
The guidance signal fundamentally changes DG-PG’s training dynamics. By reducing gradient variance (Theorem 4.2), it eliminates the need for the stabilization measures that MAPPO and IPPO require. Batch size and rollouts: DG-PG trains with mini-batches of 256–2048 and 4–12 rollouts—an order of magnitude smaller than the linear experiments—yet its actor-critic updates remain stable without split-epoch scheduling or large-batch averaging. $\alpha$ schedule: when $\alpha$ is large (early training), the strong guidance drives rapid convergence; when $\alpha$ is small (late training), the inherently stable updates allow the critic to continue learning effectively, without the actor-critic oscillation observed in MAPPO and IPPO. No return normalization: DG-PG trains stably without return normalization, which also preserves the meaningful reward differences across scenarios (e.g., 80-80 vs. 100-100) that normalization would otherwise collapse.
@@
  • High frequency (Runs 0–3; 3,000+ steps/ep): The critic learns well (explained variance reaches 0.7), but the actor undergoes entropy collapse—the policy collapses to a near-deterministic, suboptimal action within 50–70 episodes, after which performance degrades irreversibly.
@@
  • Intermediate frequency (Run 6; 300 steps/ep): The best compromise—entropy stable, critic explained variance reaching 0.35–0.49—but training rewards oscillate between 57-57 and 135-135 over 1,000 episodes with no convergence trend. Evaluation on 40 held-out scenarios yields 198-198, far worse than the Random baseline (84.2-84.2).
@@
PBRS with $\Phi(s)=-\text{Var}(\text{utilization})$ produced negligible shaping signals: $|\gamma\Phi(s^{\prime})-\Phi(s)|\approx 0.0003$ versus rewards of magnitude 6060140140. The bottleneck is gradient variance (cross-agent credit assignment), not temporal credit assignment.
@@
The dilemma is fundamentally harder for neural networks: a high-capacity critic can memorize recent scenarios but fails to generalize, while the actor’s non-convex loss landscape makes it prone to entropy collapse. Linear models ($\pi(a|s)=\text{softmax}(Ws+b)$, $V(s)=w^{\top}s+b$; ${\sim}70$ parameters) resolve this: they cannot overfit, and their convex critic loss permits more update epochs without instability. Even so, MAPPO and IPPO still require additional stabilization—split-epoch training, large mini-batches, and elevated rollout counts—as detailed in Section I. This isolates gradient variance as the sole remaining variable, providing the strictest test of DG-PG’s contribution.
@@
Each training episode consists of two phases: environment rollout (simulating $N$ agents for 3000 timesteps) and policy optimization (gradient updates on the collected experience buffer of $N\times 3000$ transitions). As $N$ grows, the optimization phase increasingly dominates wall-clock time, rising from 42% at N=5N{=}5 to 66% at N=100N{=}100, because the buffer size scales linearly with NN. All timing results were obtained on consumer-grade hardware without GPU acceleration; wall-clock times would decrease significantly with GPU-based batch gradient computation.

### v2

@@ E.1 Lemma: Guidance Consistency
$\Delta_{k+1}\leq\Delta_{k}-\frac{\eta(1-\alpha)}{2}\mathbb{E}[\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}]+\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$
By the PL condition (Assumption G.1), $\mathbb{E}[\|\nabla J(\boldsymbol{\theta}_{k})\|^{2}]\geq 2\mu\Delta_{k}$. Substituting:
$\Delta_{k+1}$ $\leq(1-(1-\alpha)\mu\eta)\Delta_{k}+\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$
Unrolling this recurrence for $T$ iterations by the geometric series formula:
$\Delta_{T}$ $\leq(1-(1-\alpha)\mu\eta)^{T}\Delta_{0}+\frac{L\eta^{2}}{2}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2})\cdot\frac{1-(1-(1-\alpha)\mu\eta)^{T}}{(1-\alpha)\mu\eta}.$
Since $\mu,\eta>0$ and $\alpha\in(0,1)$, we have $(1-\alpha)\mu\eta>0$. Combined with $\eta\leq\frac{1}{2L(1-\alpha)}$ and $\mu\leq L$ (Assumption G.1), this gives $(1-\alpha)\mu\eta\in(0,\frac{1}{2}]$. Therefore $1-(1-(1-\alpha)\mu\eta)^{T}\in(0,1]$, which gives:
$\Delta_{T}\leq(1-(1-\alpha)\mu\eta)^{T}\Delta_{0}+\frac{L\eta}{2(1-\alpha)\mu}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2}).$ (32)
To achieve $\Delta_{T}\leq\epsilon$, it suffices to bound both terms on the right-hand side of Equation 32 by $\epsilon/2$. For the second term, setting it equal to $\epsilon/2$:
$\frac{L\eta}{2(1-\alpha)\mu}(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2})=\frac{\epsilon}{2}\quad\Longrightarrow\quad\eta=\frac{(1-\alpha)\mu\epsilon}{L(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2})}.$
For the first term, requiring $(1-(1-\alpha)\mu\eta)^{T}\Delta_{0}\leq\frac{\epsilon}{2}$ gives:
$T\geq\frac{1}{(1-\alpha)\mu\eta}\log\frac{2\Delta_{0}}{\epsilon}.$
Substituting the chosen $\eta$:
$T\geq\frac{L(2\alpha^{2}B^{2}+\sigma_{\text{DG}}^{2})}{(1-\alpha)^{2}\mu^{2}\epsilon}\log\frac{2\Delta_{0}}{\epsilon}.$
@@
The quantities $L$, $\mu$, $B$, $\alpha$, and $\sigma_{\text{DG}}^{2}=\mathcal{O}(1)$ (by Theorem 4.2) are all independent of $N$. This simplifies to:
$T=\mathcal{O}\left(\frac{L}{\mu^{2}\epsilon}\log\frac{1}{\epsilon}\right),$
completing the proof. ∎
@@ H.1 Physical Environment
The efficiency parameters $\eta_{\text{CPU}}$ and $\eta_{\text{Mem}}$ model performance-per-watt differences across hardware generations. Gen-1 instances (circa 2015) have $\eta_{\text{CPU}}\in[0.70,0.75]$, Gen-2 (circa 2017) have $\eta_{\text{CPU}}\in[0.84,0.98]$, and Gen-3 (circa 2020+) have $\eta_{\text{CPU}}\in[1.06,1.08]$. Job completion time scales as $\text{duration}/\eta_{\text{CPU}}$, making high-efficiency servers more productive. Server types are sampled according to the specified weights and randomly shuffled across indices at scenario generation to prevent agents from exploiting positional patterns.
@@ H.3 Reward Function
where $Q_{\text{global}}(t)$ is the global buffer size, $Q_{k}(t)$ is the local queue at server $k$, $L_{k}(t)$ is the workload (CPU + memory) at server $k$, $\eta_{k}$ is the server’s efficiency, and $C_{\text{total}}$ is the total cluster capacity. We set $w_{1}=1$ and $w_{2}=20$ so that both penalties contribute comparably to the reward. Without the multiplier, the queue penalty typically ranges in $[0,20]$, whereas the normalized energy ratio remains below $1$. Load-balancing (utilization variance) is handled entirely by the DG-PG guidance signal and is therefore excluded from the reward.
@@ I.2 Linear Models (Controlled Comparison)
  • Large mini-batch (10,000) and high rollout count (24): With smaller batches (e.g., 256–1,024), MAPPO and IPPO exhibit wildly oscillating rewards that never converge. Multi-agent policy gradients in this environment have high variance due to heterogeneous servers, bimodal workloads, and non-stationary arrivals, requiring large batches to average over noise.
@@ I.3 Neural Network Models (DG-PG)
  • Guidance weight $\alpha$ follows a dynamic three-phase schedule (Section 6.2): $\alpha=0.9$ for the first 10% of training, linear decay $0.9\to 0.2$ from 10% to 50%, and $\alpha=0.2$ thereafter
@@
The guidance signal fundamentally changes DG-PG’s training dynamics. By reducing gradient variance (Theorem 4.2), it eliminates the need for the stabilization measures that MAPPO and IPPO require. Batch size and rollouts: DG-PG trains with mini-batches of 256–2048 and 4–12 rollouts, an order of magnitude smaller than the linear experiments, yet its actor-critic updates remain stable without split-epoch scheduling or large-batch averaging. $\alpha$ schedule: when $\alpha$ is large (early training), the strong guidance drives rapid convergence; when $\alpha$ is small (late training), the inherently stable updates allow the critic to continue learning effectively, without the actor-critic oscillation observed in MAPPO and IPPO. No return normalization: DG-PG trains stably without return normalization. This preserves meaningful reward differences across scenarios (e.g., 80-80 vs. 100-100) that normalization would otherwise collapse.
@@
  • High frequency (Runs 0–3; 3,000+ steps/ep): The critic learns well (explained variance reaches 0.7), but the actor undergoes entropy collapse, converging to a near-deterministic, suboptimal action within 50–70 episodes, after which performance degrades irreversibly.
@@
  • Intermediate frequency (Run 6; 300 steps/ep): The best compromise, with entropy stable and critic explained variance reaching 0.35–0.49, but training rewards oscillate between 57-57 and 135-135 over 1,000 episodes with no convergence trend. Evaluation on 40 held-out scenarios yields 198-198, far worse than the Random baseline (84.2-84.2).
@@
PBRS with $\Phi(s)=-\text{Var}(\text{utilization})$ produced negligible shaping signals, with $|\gamma\Phi(s^{\prime})-\Phi(s)|\approx 0.0003$ versus rewards of magnitude 6060140140. The bottleneck is gradient variance (cross-agent credit assignment), not temporal credit assignment.
@@
The dilemma is fundamentally harder for neural networks. A high-capacity critic can memorize recent scenarios but fails to generalize, while the actor’s non-convex loss landscape makes it prone to entropy collapse. Linear models ($\pi(a|s)=\text{softmax}(Ws+b)$, $V(s)=w^{\top}s+b$; ${\sim}70$ parameters) resolve this because they cannot overfit, and their convex critic loss permits more update epochs without instability. Even so, MAPPO and IPPO still require additional stabilization, including split-epoch training, large mini-batches, and elevated rollout counts, as detailed in Section I. This isolates gradient variance as the sole remaining variable, providing the strictest test of DG-PG’s contribution.
@@
Each training episode has two phases: environment rollout (simulating $N$ agents for 3000 timesteps) and policy optimization (gradient updates on the collected experience buffer of $N\times 3000$ transitions). As $N$ grows, the optimization phase increasingly dominates wall-clock time, rising from 42% at N=5N{=}5 to 66% at N=100N{=}100, because the buffer size scales linearly with NN. All timing results were obtained on consumer-grade hardware without GPU acceleration. Wall-clock times would decrease significantly with GPU-based batch gradient computation.